OpenStudy (anonymous):

Precal. Help. Please. log(x) + log(x+2) = 3

OpenStudy (dumbcow):

logx +logy = logxy use this property log(x(x+2)) = 3 put it in exponential form x(x+2) = 10^3 set equal to 0 x^2 +2x - 1000 =0 use quadratic formula \[\frac{-2\pm \sqrt{4004}}{2}\]

OpenStudy (anonymous):

I did all of that up to the quadratic formula! Thank you!

OpenStudy (dumbcow):

you can reduce that too -1 +- sqrt(1001)

OpenStudy (anonymous):

the root -1-sqrt(1001) is not acceptable...only +ve values should be taken in the domain of log